OpenStudy (anonymous):

HELP PLEASE! How do i solve this? step by step f(x)=2^x-1 g(x)=sqrt x

OpenStudy (dumbcow):

what are we solving? you just gave 2 functions?

OpenStudy (anonymous):

I have to find the approximate solutions. then i have to graph it.

OpenStudy (anonymous):

i can do the graphing

OpenStudy (anonymous):

wait are the vertices the approx. values?

OpenStudy (dumbcow):

i dont understand, to solve an equation you need something on left side and right side 2^x-1 = ? sqrt(x) = ?

OpenStudy (anonymous):

Page 3 problem 1

OpenStudy (anonymous):

Unless i have to somehow put tem together and use the quadratic equation

OpenStudy (dumbcow):

the worksheet doesn't really explain anything but if these are meant to be supply/demand functions then maybe the solutions are supposed to be the equilibrium point f(x) = g(x) 2^x -1 = sqrt(x) if you graph both functions you will see that they both meet at the origin and at 1 other point use graphing software to estimate the point you cant solve this using algebra

OpenStudy (anonymous):

Oh Okay Well Thank You Anyway! And Yeah I Graphed Them But Didnt think that could be it.

OpenStudy (triciaal):

the picture has f(x) = 2^x your post has f(x)= 2^x-1

OpenStudy (anonymous):

No thats the very first problem which ive already done i mean the first problem under solving systems graphicaly on page 3 look again :)

OpenStudy (anonymous):

Thank You So Much!

OpenStudy (triciaal):

I see the other problem as well did you get the solution?

OpenStudy (anonymous):

Yes yes i did :)